# Identifying Halos in Cosmological Simulations with Continuous Wavelet Analysis: The 2D Case

### v2

@@ 6 Conclusions and Discussions
In this paper, we employ a wavelet-based algorithm to identify halos within pseudo-2D data generated from the SIMBA simulation and compare the results with the traditional FOF method. Our algorithm leverages the ability to discern structures across both space and scale via the CWT, which can be considered as a ‘mathematical microscope’ for delving into the intricate structure of the density field and automatically extracting its clustering properties. To locate the halos in position-scale space, we initially select a range of scale parameters $w$ and compute the grid CWT for a series of $w_{i}$ within the selected range. We then identify the local maxima of the CWT at each scale and compare them across scales to retain the most prominent maximum at each position. Once all local maxima have been detected, we segment the CWT grid into distinct ‘islands’ to delineate the shape of these local maxima. The ‘islands’ are zoomed and projected back into the simulation space from large scale to small scale, discarding any overlapping islands originating from larger scales. Subsequently, each particle is assessed to ascertain whether it is part of an island. Halo particles are designated, and only halos comprising more than 5 particles are retained.
To prevent the incorrect detection of Poisson noise as halos, we conduct Monte Carlo simulations to assess the significance of each local maximum. We randomly introduce mock particles, equivalent in number to the real data, into the simulation space and determine the detection threshold and its corresponding significance level from the histogram of the CWT values of the mock data. To eliminate the time-consuming aspect of Monte Carlo calculations, we pre-run simulations across a broad range of particle numbers and scales. The outcomes are then fitted to an empirical formula, enabling the establishment of a threshold at the 99.5% significance level with an error margin of approximately 10%. Additionally, the impact of resolution is examined. If the resolution is too low, it becomes challenging to differentiate between closely situated halos, leading to a reduction in the number of detected halos (to two-thirds of the standard resolution). Conversely, doubling the resolution yields a more refined halo boundary, yet it may not accurately reflect the true halo structure because of the shift of its maxima, even though it requires four times the computational resources.
@@
    Based on the halo mass function, it can be inferred that in the lowest mass range (around 5 particles), the CWT halo exhibits the highest excess in the number of halos, resulting in a finer shape of its mass function. After the bulge of the FOF halo mass function, the number of CWT halos gradually increases to match that of the FOF and ultimately stabilizes at an excess of approximately 20%.
@@
    Despite having a lower amplitude, the halo power spectrum of the CWT halos is consistent with that of the FOF halos across all scales. This suggests that, although the clustering intensity of the CWT halos is weaker than that of the FOF halos, the trends they follow are similar.
Overall, our findings indicate that the CWT is capable of identifying halos in cosmological simulations, owing to its high consistency with the identification results of the FOF method. The aim of this paper is to demonstrate the feasibility of our CWT approach. It should be noted that there is significant potential for improvement in both our program and algorithm, as discussed in Appendix A. Additionally, our algorithm relies solely on the positional data of these particles. To ascertain that any group of particles constitutes a halo that more closely resembles reality, we require the complete dynamic information of the particles. This is because dark matter halos are gravity-bound, virialized structures. Consequently, in future work, we will incorporate the velocity information of particles to achieve a more accurate representation of the dynamical structure of our CWT-identified halos.
In the current work, we chose to work with 2D rather than 3D data for several reasons as follows. (1) Compared with the 3D case, halos identified with 2D data can be completed with lower costs in processing time, memory and storage space, making the development and testing of algorithms more efficient. (2) Working in the 2D case can simplify the problem, allowing us to more easily identify and address issues without the added complexity of an additional dimension. (3) 2D data is easier to visualize, which is very helpful for demonstrating and understanding the effects of an algorithm. (4) Working in the 2D case can serve as an intermediate step towards transitioning to the 3D case; the concept can first be proven in 2D before extending the research to 3D case. Despite these benefits, our 2D test is not yet practical for real-world applications, given that we inhabit a 3D universe. To develop a program capable of analyzing real data rather than pseudo-2D data, we must account for the 3D case and optimize our program for enhanced performance.
Currently, the performance of our program results in unacceptable time costs when processing significantly larger datasets in the 3D context, so we particularly need a fast wavelet algorithm for 3D particle data (e.g. Romeo et al., 2008; Wang & He, 2023), which will be the target of our next study. Furthermore, the capability to discern structures across scales using CWT holds the potential to detect substructures within the CWT halo, which may also be explored in our future research.
@@ Acknowledgments
MXL thanks Prof. Pin Lyu for his assistance with some numerical techniques. The authors also express their gratitude to the SIMBA team for making their data publicly available.
@@ Appendix A Performance Evaluation of the CWT Method
Figure 13: The computation time varies with different data sizes, ranging from 66,942 to 1,435,111 (which corresponds to a thickness from $0.02\,{h^{-1}{\rm Mpc}}$ to $0.5\,{h^{-1}{\rm Mpc}}$). The blue dashed line represents the computation time for the FOF method, and the solid line represents that for the CWT method. The $\mathcal{O}(N\log N)$ time complexity for FOF method exhibits as linear in this log-log plot.
To properly evaluate the performance of an algorithm, it is crucial to understand its actual capabilities. In the case of our CWT method, we conducted tests to measure the time required for calculating data of varying sizes. We select slices with thicknesses ranging from $0.02\,{h^{-1}{\rm Mpc}}$ to $0.5\,{h^{-1}{\rm Mpc}}$ to generate test pseudo-2D data, which we then use to run our halo identification program. The tests are conducted on a platform equipped with an AMD Ryzen 9 7900X CPU and 64GB of DDR5 6000MHz RAM.
The program consists of two components: calculating grid CWT and characterizing local maxima. We test the time costs for each part separately and compare them with those of the FOF method. The results are depicted in Figure 13. Both the FOF method, as calculated by yt-project, and our CWT method exhibit linear time complexity in our tests. The time cost for the former part of our program (represented by the red line) increases linearly with data size, while the latter part demonstrates significantly less variation. This is because, even with fewer particles in the thinnest slice, the number of local maxima does not change rapidly. The slice at $0.02\,{h^{-1}{\rm Mpc}}$ contains 2,481 local maxima, and concurrently, the thickest slice at $0.5\,{h^{-1}{\rm Mpc}}$ contains only 4,270, hence the time required to characterize them varies correspondingly slowly.
The performance of our program is significantly inferior to that of YT. We spend 60 times as much time on the smallest data set and 40 times as much on the largest data set. However, it is not advisable to compare the performance of a test program directly with that of a well-optimized project. Therefore, our primary goal is to confirm the $\mathcal{O}(N)$ time complexity of our program.

### v3

@@ 6 Conclusions and Discussions
In this paper, we employ a wavelet-based algorithm to identify halos within the pseudo-2D data generated from the SIMBA simulation and compare the results with the traditional FOF method. Our algorithm leverages the ability to discern structures across both space and scale via the CWT, which can be considered as a ‘mathematical microscope’ for delving into the intricate structure of the density field and automatically extracting its clustering properties. To locate the halos in position-scale space, we initially select a range of scale parameters $w$ and compute the grid CWT for a series of $w_{i}$ within this range. We then identify the local maxima of the CWT at each scale and compare them across scales to retain the most prominent maximum at each position. Once all local maxima have been detected, we segment the CWT grid into distinct islands to delineate the shape of these local maxima. The islands are zoomed and projected back into the simulation space in descending order of scale, discarding any overlapping islands originating from smaller scales. Subsequently, each particle is assessed to ascertain whether it is part of an island. Halo particles are designated, and only halos comprising more than five particles are retained.
To prevent the incorrect detection of Poisson noise as halos, we conduct Monte Carlo simulations to assess the significance of each local maximum. By introducing a random set of mock particles, equal in number to the actual data, into the simulation space, we can establish a detection threshold and its significance level. This is done by analyzing the histogram of the CWT values obtained from the mock data. To eliminate the time-consuming aspect of Monte Carlo calculations, we prerun simulations across a broad range of particle numbers and scales. The outcomes are then fitted to an empirical formula. This formula allows us to set a threshold at the 99.5% significance level with an error margin of approximately 10%.
Furthermore, we examine the effect of resolution on halo detection. Low resolution can make it difficult to distinguish between closely located halos, potentially reducing the number of detected halos to two-thirds of what would be found at standard resolution. On the other hand, doubling the resolution provides a more detailed definition of halo boundaries. However, this increased resolution might not accurately represent the true halo structure due to shifts in the location of maxima, and it demands a computational cost four times greater than that of the standard resolution.
@@
    Based on the cumulative halo mass function, it can be inferred that in the lowest mass range (around five particles), the CWT halo shows the greatest excess in the number of halos and then drops sharply around 10 particles. As the halo mass grows larger, the number of CWT halos gradually increases to match that of the FOF, and ultimately stabilizes at an excess of approximately 20%.
@@
    Despite having a lower amplitude, the halo power spectrum of the CWT halos is consistent with that of the FOF halos across all scales. This indicates that while the clustering intensity of CWT halos may be less pronounced than that of FOF halos, the overall trends they follow are similar.
Overall, our findings indicate that the CWT is capable of identifying halos in cosmological simulations, owing to its high consistency with the FOF method. The aim of this paper is to demonstrate the feasibility of our CWT approach. It should be noted that there is significant potential for improvement in both our program and algorithm, as discussed in the Appendix A. Additionally, our algorithm relies solely on the positional data of these particles. To ascertain that any group of particles constitutes a halo that more closely resembles reality, we require the complete dynamic information of the particles. This is because dark matter halos are gravity-bound, virialized structures. Consequently, in future work, we will incorporate the velocity information of particles to achieve a more accurate representation of the dynamical structure of our CWT-identified halos.
In the current work, we chose to work with 2D rather than 3D data for several reasons as follows. (1) Compared with the 3D case, halos identified with 2D data can be completed with lower costs in processing time, memory, and storage space, making the development and testing of algorithms more efficient. (2) Working in the 2D case can simplify the problem, allowing us to more easily identify and address issues without the added complexity of an additional dimension. (3) 2D data is easier to visualize, which is very helpful for demonstrating and understanding the effects of an algorithm. (4) Working in the 2D case can serve as an intermediate step toward transitioning to the 3D case; the concept can first be proven in 2D before extending the research to the 3D case. Despite these benefits, our 2D test is not yet practical for real-world applications, given that we inhabit a 3D universe. To develop a program capable of analyzing real data rather than pseudo-2D data, we must account for the 3D case and optimize our program for enhanced performance.
Currently, the performance of our program incurs unacceptable time costs when processing significantly larger 3D datasets, necessitating a fast wavelet algorithm for 3D particle data (e.g., Romeo et al., 2008; Wang & He, 2023), which will be the target of our next study. Furthermore, the capability to discern structures across scales using CWT holds the potential to detect substructures within CWT halos. This capability presents an exciting avenue for exploration in our future research.
@@ Acknowledgments
The authors thank the anonymous referee for helpful comments and suggestions. M.X.L. thanks Prof. Pin Lyu for his assistance with some numerical techniques. The authors also express their gratitude to the SIMBA team for making their data publicly available.
@@ Appendix A Performance Evaluation of the CWT Method
Figure 13: The computation time varies with different data sizes, ranging from 66,942 to 1,435,111 (which corresponds to a thickness from $0.02\,{h^{-1}{\rm Mpc}}$ to $0.5\,{h^{-1}{\rm Mpc}}$). The blue dashed line represents the computation time for the FOF method, while the solid line represents that for the CWT method, both measured in seconds. The $\mathcal{O}(N\log N)$ time complexity for the FOF method is exhibited as linear in this log-log plot.
To properly evaluate the performance of an algorithm, it is crucial to understand its actual capabilities. For our CWT method, we conducted tests to measure the time required for calculating data of varying sizes. We select slices with thicknesses ranging from $0.02\,{h^{-1}{\rm Mpc}}$ to $0.5\,{h^{-1}{\rm Mpc}}$ to generate test pseudo-2D data for our halo identification program. The tests are conducted on a platform equipped with an AMD Ryzen 9 7900X CPU and 64GB of DDR5 6000 MHz RAM.
The program consists of two components: calculating grid CWT and characterizing local maxima. We test the time costs for each part individually and compare them with those of the FOF method. The results are depicted in Figure 13. Both the FOF method, as calculated by the yt project, and our CWT method exhibit linear time complexity in our tests. The time cost for the former part of our program (represented by the red line) increases linearly with data size, while the latter part demonstrates significantly less variation. This is because, even with fewer particles in the thinnest slice, the number of local maxima does not change rapidly. The slice of $0.02\,{h^{-1}{\rm Mpc}}$ contains 2,481 local maxima, and concurrently, the thickest slice of $0.5\,{h^{-1}{\rm Mpc}}$ contains only 4270; hence, the time required to characterize them varies correspondingly slowly.
The performance of our program is significantly inferior to that of the yt project. We spend 60 times more computation time on the smallest data set and 40 times more on the largest data set. However, it is not advisable to compare the performance of a test program directly with that of a well-optimized project. Therefore, our primary goal is to confirm the $\mathcal{O}(N)$ time complexity of our program.